(5x-6p)^2+(7z-5x)^2+(6p-7z)^2Factorised this equation

Grade:8

1 Answers

Anish Singhal
5 years ago
Take 5x-6p=a, 7z-5x=b, 6p-7z=c and then use this formula
(a+b+c)2–2(ab+ac+bc)=(a2+b2+c2)